# Decoupled Training: Return of Frustratingly Easy Multi-Domain Learning

### v1

@@ Abstract
Multi-domain learning (MDL) aims to train a model with minimal average risk across multiple overlapping but non-identical domains. To tackle the challenges of dataset bias and domain domination, numerous MDL approaches have been proposed from the perspectives of seeking commonalities by aligning distributions to reduce domain gap or reserving differences by implementing domain-specific towers, gates, and even experts. MDL models are becoming more and more complex with sophisticated network architectures or loss functions, introducing extra parameters and enlarging computation costs. In this paper, we propose a frustratingly easy and hyperparameter-free multi-domain learning method named Decoupled Training (D-Train). D-Train is a tri-phase general-to-specific training strategy that first pre-trains on all domains to warm up a root model, then post-trains on each domain by splitting into multi-heads, and finally fine-tunes the heads by fixing the backbone, enabling decouple training to achieve domain independence. Despite its extraordinary simplicity and efficiency, D-Train performs remarkably well in extensive evaluations of various datasets from standard benchmarks to applications of satellite imagery and recommender system.
@@ Introduction
The success of deep learning models across a wide range of fields often relies on the fundamental assumption that the data points are independent and identically distributed (i.i.d.). However, in real-world scenarios, training and test data are usually collected from different regions, devices, or platforms, consisting of multiple overlapping but non-identical domains. For example, a popular satellite dataset named Functional Map of the World (FMoW)  (Christie et al. 2018), which aims to predict the functional purpose of buildings and land use on this planet, contains large-scale satellite images from various regions with different appearances and styles. In this case, jointly training a single model obscures domain distinctions, while separately training multiple models by domains reduces training data in each model (Joshi et al. 2012). This dilemma motivated the research on multi-domain learning (Joshi et al. 2012; Liu, Qiu, and Huang 2017; Ma et al. 2018; Alice et al. 2019; Tang et al. 2020). To figure out the challenges of multi-domain learning, we first delved into these two standard benchmark datasets. By analyzing the examples of these datasets, it is obvious that dataset bias across domains is one of the biggest obstacles to multi-domain learning. As shown in Figure 1(a), reviews from different domains have different keywords and styles. Further, Figure 1(b) reveals that images of Office-Home are from four significantly different domains: Art, Clipart, Product and Real-World, with various appearances and backgrounds. To tackle the dataset bias problem, numerous approaches have been proposed and they can be briefly grouped into two categories: 1) Seeking Commonalities. Some classical solutions (Liu, Qiu, and Huang 2017; Alice et al. 2019) adopting the insightful idea of domain adversarial training have been proposed to extract domain-invariant representations across multiple domains. 2) Reserving Differences. These approaches  (Ma et al. 2018; Tang et al. 2020) adopt multi-branch network architectures with domain-specific towers, gates, and even experts, implementing domain-specific parameters to avoid domain conflict caused by dataset bias across domains. As shown in Figure 2, these methods are becoming more and more complex with sophisticated network architectures or loss functions.
@@
Figure 1: (a)-(b): Review examples from a recommender system benchmark named Amazon with various styles and keywords, as well as visual examples from a computer vision dataset named Office-Home with various appearances and backgrounds, reveal the main challenge of dataset bias. (c)-(d): The distribution of sample number across domains is naturally imbalanced or even long-tailed, indicating another major challenge of domain domination.
Further, different from multi-task learning which focuses on tackling different tasks within a single domain, multi-domain learning shares the same label space across multiple domains with different marginal distributions. We thus calculated the distribution of sample number across domains on several benchmark datasets and the results are shown in Figure 1. In practice, the distributions of sample number across domains are usually twisted, imbalanced, or even long-tailed, causing another main challenge of multi-domain learning: domain domination. In this case, some tailed domains may have much fewer examples and it would be difficult to train satisfied models on them. Without proper network designs or loss functions, the model will be easily dominated by the head domains with much more examples and shift away from the tailed ones with few examples. Realizing the main challenges of dataset bias and domain domination in multi-domain learning, we aim at proposing a general, effective, and efficient method to tackle these obstacles all at once. We first rethought the development of multi-domain learning and found that the approaches in this field are becoming more and more sophisticated, consisting of multifarious network architectures or complex loss functions with many trade-off hyperparameters. By assigning different parameters across domains, these designs may be beneficial in some cases but they will absolutely include more parameters and enlarge computation cost, as well as introduce much more hyperparameters. For example, the latest state-of-the-art method named PLE has to tune the numbers of domain-specific experts and domain-agnostic experts in each layer and design the network structures of each expert, each tower, and each gate network. Motivated by the famous quote of Albert Einstein, “everything should be made as simple as possible, but no simpler”, we proposed a frustratingly easy multi-domain learning method named Decoupled Training (D-Train). D-Train is a training strategy based on the original frustratingly easy but most general shared-bottom architecture. D-Train is a tri-phase general-to-specific training strategy that first pre-trains on all domains to warm up a root model, then post-trains on each domain by splitting into multi-heads, and finally fine-tunes the heads by fixing the backbone, enabling decouple training to achieve domain independence. Despite its extraordinary simplicity and efficiency, D-Train performs remarkably well in extensive evaluations of various datasets from standard benchmarks to applications of satellite imagery and recommender system. In summary, this paper has the following contributions:
@@
To tackle the dataset bias problem in multi-domain learning, various approaches (Liu, Qiu, and Huang 2017; Alice et al. 2019) have been proposed from the perspective of domain alignment, by adopting the insightful idea of domain adversarial training to extract domain-invariant representations across domains. To smooth the presentation of domain alignment in MDL, we will first give a brief review of domain adaptation. There are mainly two categories of domain adaptation formulas: covariate shift (Quionero-Candela et al. 2009; Pan and Yang 2010; Long et al. 2015; Ganin and Lempitsky 2015) and label shift (Lipton, Wang, and Smola 2018; Azizzadenesheli et al. 2019; Alexandari, Kundaje, and Shrikumar 2020), while we focus on the former in this paper since it is more relevant with the topic of MDL. Recent deep domain adaptation methods tackle domain shifts from the perspectives of either moment matching or adversarial training, in which the former aligns feature distributions by minimizing the distribution discrepancy across domains (Long et al. 2015; Tzeng et al. 2014; Long et al. 2017). Further, domain adversarial neural network (DANN) (Ganin et al. 2016) becomes the mainstream method in domain adaptation. It introduces a domain discriminator to distinguish the source features from the target ones, while the feature extractor is designed to confuse the domain discriminator. In this way, the domain discriminator and feature extractor are competing in a two-player minimax game. Its natural extension to MDL is DANN-MDL. Later, CDAN (Long et al. 2018) further tailors the discriminative information conveyed in the classifier predictions into the input of the domain discriminator, whose natural extension to MDL is CDAN-MDL. Following the main idea of the minimax game, several variants of adversarial training methods (Pei et al. 2018; Tzeng et al. 2017; Saito et al. 2018; Zhang et al. 2019) were proposed. MulANN (Alice et al. 2019) tailors the insight of domain adversarial training into the MDL problem by introducing a domain discriminator into the shared model with a single head. In contrast, ASP-MTL (Liu, Qiu, and Huang 2017) includes a shared-private model and a domain discriminator.
@@
Another series of methods (Ma et al. 2018; Tang et al. 2020; Sheng et al. 2021) for multi-domain learning adopt multi-branch network architectures and develop domain-specific parameters (Dredze, Kulesza, and Crammer 2010) to avoid domain conflict caused by dataset bias across domains. Among them, Shared Bottom (SB) (Ruder 2017) is the frustratingly easy but effective one. Further, MoE (Jacobs et al. 1991) and its extension of MMoE (Ma et al. 2018) adopt the insightful idea of the mixture of experts to learn different mixture patterns of experts assembling, respectively. PLE (Tang et al. 2020) explicitly separates domain-shared and domain-specific experts to alleviate harmful parameter interference across domains. Note that, PLE further applies progressive separation routing with several deeper layers but we only adopt one layer for a fair comparison with other baselines. Other MDL methods focus on maintaining shared and domain-specific parameters by confidence-weighted combination (Dredze, Kulesza, and Crammer 2010), domain-guided dropout (Xiao et al. 2016), or prior knowledge about domain semantic relationships (Yang and Hospedales 2015). Meanwhile, various task-specific MDL approaches have been proposed for computer vision (Rebuffi, Bilen, and Vedaldi 2018; Mancini et al. 2020; Nam and Han 2016; Rebuffi, Bilen, and Vedaldi 2017; Li and Vasconcelos 2019; Fourure et al. 2017), natural language processing (Wu and Guo 2020; Williams 2013; Pham, Crego, and Yvon 2021) and recommender system (Hao et al. 2021; Chen et al. 2020; Du et al. 2019; Gu et al. 2021; Li et al. 2021b, a; Salah, Tran, and Lauw 2021). The comparison between these MDL methods are summarized in Figure 2.
@@
This paper aims at proposing a simple, effective, and frustratingly easy method for multi-domain learning (MDL). Given data points from multiple domains $\{\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{T}\}$, $T$ is the domain number. Denote $\mathcal{D}_{t}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n_{t}}$ the data from domain $t$, where $\mathbf{x}_{i}$ is an example, $\mathbf{y}_{i}$ is the associated label and $n_{t}$ is the sample number of domain $t$. Denote the shared backbone $\psi$ and the domain-specific heads $\{{h}_{1},{h}_{2},...,{h}_{T}\}$ respectively. The goal of MDL is to improve the performance of each domain $\mathcal{D}_{t}$. As mentioned above, the proposed Decoupled Training (D-Train) is a frustratingly easy multi-domain learning method based on the original shared-bottom architecture. D-Train takes a general-to-specific training strategy. It first pre-trains on all domains to warm up a root model. Then, it post-trains on each domain by splitting into multi-heads. Finally, it fine-tunes the heads by fixing the backbone, enabling decouple training to achieve domain independence.
@@
As mentioned before, jointly training a single model obscures domain distinctions, leading to domain conflict caused by the specificity of different domains. To reflect the domain specificity and achieve satisfactory performance for each domain, we adopt the shared-bottom architecture that has a shared feature extractor $\psi$ and various domain-specific heads $\{{h}_{1},{h}_{2},...,{h}_{T}\}$ to tackle the challenge of dataset bias across domains. With this design, the parameters of the feature extractor will be updated simultaneously by gradients of samples from all domains, but the parameters of domain-specific heads are trained on each domain. The optimization function of the post-training phase over all domains can be formalized as
@@
(a) Domain Domination
(b) After Post-train
(c) After Fine-tune
@@
Fine-tune: decouple-train for domain independence
Regarding the benefits of domain-specific parameters across domains, a natural question arises: Can the shared-bottom architecture fully solve the problem of domain domination? To answer this question, we adopt the Euclidean norm to calculate the parameter update between the domain-specific heads $\{{h}_{1},{h}_{2},...,{h}_{T}\}$ and the domain-agnostic head ${h}_{0}$ at the pre-training phase on DomainNet. As shown in Figure 4(a), the domain of Real has much more examples than other domains, and is believed to dominate the training process. However, as shown in Figure 4(b), the parameter update after the post-train phase reveal that the training process is still dominated by the head domain which still has the largest value of the parameter update. Hence, though the challenge of dataset bias across domains can be alleviated by introducing domain-specific heads in the shared-bottom architecture, the domain domination problem still exists after the post-train phase. This is reasonable since the domain-shared parameters of the backbone will be dominated by the head domains. To this end, we propose a decoupling training strategy by fully fixing the parameters of the feature extractor to achieve domain independence. Formally,
In this way, the parameters of the domain-specific heads will be learned by samples from each domain as shown in Figure 3(a). With this kind of domain-independent training, the head domains will no longer dominate the training of the tailed domains at this phase. Further, the parameter update between phases becomes more balanced across domains as shown in Figure 4(c). Meanwhile, the training curves as shown in Figure 3(b) reveal that fine-tuning across domains can further improve the performance.
Why does D-Train work?
As mentioned before, dataset bias and domain domination are the main challenges of multi-domain learning. All of the existing works, including the methods from the perspectives of both seeking commonalities or reserving differences, have to face the challenge of caring for this and losing that. With the shared parameters across domains, these methods will influence each other. When the problem of domain domination or the domain conflict caused by dataset bias cannot be ignored, the MDL model will struggle in finding an optimal solution for all domains. Actually, as shown in the training process of DomainNet in Figure 3(b), different domains achieve the optimal performance at different time stamps after the post-training phase. However, the goal of multi-domain learning is to use one model to serve all domains, and thus we cannot use these models simultaneously. On the contrary, with the proposed decoupling training strategy, different domains in D-Train will train independently at the fine-tuning phase and reach an optimal solution for each domain.
@@
In this section, we compared the proposed D-Train method with three categories of baselines: Single Branch, Domain Alignment, and Multiple Branch.
@@
In this section, we adopt two standard benchmarks in the field of domain adaptation with various dataset scales, in which Office-Home is in a low-data regime and DomainNet is a large-scale one. D-Train and all baselines in this section are implemented in a popular open-sourced library  [^1] https://github.com/thuml/Transfer-Learning-Library which has implemented a high-quality code base for many domain adaptation baselines.
@@
Office-Home is a standard multi-domain learning dataset  (Venkateswara et al. 2017) with $65$ classes and $15,500$ images from four significantly different domains: Art, Clipart, Product, and Real-World. As shown in Figure 1(b), there exist challenges of dataset bias and domain domination in this dataset. Following existing works on this dataset, we adopt ResNet-50 as the backbone and randomly initialize fully connected layers as heads. We set the learning rate as $0.0003$ and batch size as $24$ in each domain for D-Train and all baselines. As shown in Table 1, Separately Training is a strong baseline and even outperforms Jointly Training, since the latter obscures domain distinctions and cannot tackle the dataset bias across domains. Multi-Domain Learning methods from the perspective of domain alignment work much better by introducing a domain discriminator and exploiting the domain information. However, applying domain alignment is not an optimal solution in MDL since the domain gap can only be reduced but not removed. Finally, the proposed D-Train consistently improves on all domains, even the tailed domain of “Art”. D-Train achieves a new state-of-the-art result with an average accuracy of $86.0$.
@@
DomainNet (Peng et al. 2019) is a large-scale multi-domain learning and domain adaptation dataset with $345$ categories. We utilize $4$ domains with different appearances including Clipart, Painting, Real, and Sketch where each domain has about $40,000$ to $200,000$ images. As shown in Figure 4(a), the domain of ”Real” has much more examples than other domains, and is believed to dominate the training process. Following the code base in Transfer Learning Library, we adopt mini-batch SGD with the momentum of $0.9$ as an optimizer, and the initial learning rate is set as $0.01$ with an annealing strategy. We adopt ResNet-101 as the backbone since DomainNet is much larger and more difficult than the previous Office-Home dataset. Meanwhile, the batch size is set as $20$ in each domain here for D-Train and all baselines. As shown in Table 2, the proposed D-Train outperforms all baselines, no matter measured by average accuracy or worst accuracy in all domains. Since the domain of “Real” has much more samples than other domains, it achieves competitive performance while separately training. However, other domains benefit from training via D-Train.
@@
In this section, we adopt a popular satellite dataset named Functional Map of the World (FMoW)  (Christie et al. 2018), which aims to predict the functional purpose of buildings and land use on this planet, contains large-scale satellite images with different appearances and styles from various regions: Africa, Americas, Asia, Europe, and Oceania. FMoW is a natural dataset for multi-domain learning. D-Train and all baselines in this section are implemented in a popular open-sourced library named WILDS  [^2] https://github.com/p-lambda/wilds since it enables easy manipulation of this dataset. Each input $\mathbf{x}$ in FMoW is an RGB satellite image that is resized to $224\times 224$ pixels and the label $y$ is one of $62$ building or land use categories. For all experiments, we follow (Christie et al. 2018) and use a DenseNet-121 model (Huang et al. 2017) pretrained on ImageNet. We set the batch size to be $64$ on all domains. Following WILDS, we report the average accuracy and worst-region accuracy in all multi-domain learning methods.
@@
As shown in Table 3, it’s not wise to train a separate model for each domain on FMoW, since the data on some domains is extremely scarce. Joint Train or Shared Bottom is also not optimal, because conflicts widely exist in some domains, such as the Europe domain and Africa domain which have very different appearances. Note that D-Train also outperforms all baselines in this difficult dataset.
Table 4: AUC (%) on Amazon Product Review for multi-domain learning.

### v2

@@ Abstract
Multi-domain learning (MDL) aims to train a model with minimal average risk across multiple overlapping but non-identical domains. To tackle the challenges of dataset bias and domain domination, numerous MDL approaches have been proposed from the perspectives of seeking commonalities by aligning distributions to reduce domain gap or reserving differences by implementing domain-specific towers, gates, and even experts. MDL models are becoming more and more complex with sophisticated network architectures or loss functions, introducing extra parameters and enlarging computation costs. In this paper, we propose a frustratingly easy and hyperparameter-free multi-domain learning method named Decoupled Training (D-Train). D-Train is a tri-phase general-to-specific training strategy that first pre-trains on all domains to warm up a root model, then post-trains on each domain by splitting into multi-heads, and finally fine-tunes the heads by fixing the backbone, enabling decouple training to achieve domain independence. Despite its extraordinary simplicity and efficiency, D-Train performs remarkably well in extensive evaluations of various datasets from standard benchmarks to applications of satellite imagery and recommender systems.
@@ Introduction
The success of deep learning models across a wide range of fields often relies on the fundamental assumption that the data points are independent and identically distributed (i.i.d.). However, in real-world scenarios, training and test data are usually collected from different regions or platforms, consisting of multiple overlapping but non-identical domains. For example, a popular satellite dataset named FMoW  (Christie et al. 2018), which predicts the functional purpose of buildings and land use on this planet, contains large-scale satellite images from various regions with different appearances and styles. Thus, jointly training a single model obscures domain distinctions, while separately training multiple models by domains reduces training data in each model (Joshi et al. 2012). This dilemma motivated the research on multi-domain learning (Joshi et al. 2012; Liu, Qiu, and Huang 2017; Ma et al. 2018; Alice et al. 2019; Tang et al. 2020).
To figure out the challenges of multi-domain learning, we first delved into these two standard benchmark datasets. By analyzing the examples of these datasets, it is obvious that dataset bias across domains is one of the biggest obstacles to multi-domain learning. As shown in Figure 1(a), reviews from different domains have different keywords and styles. Further, Figure 1(b) reveals that images of Office-Home are from four significantly different domains: Art, Clipart, Product and Real-World, with various appearances and backgrounds. To tackle the dataset bias problem, numerous approaches have been proposed and they can be briefly grouped into two categories: 1) Seeking Commonalities. Some classical solutions (Liu, Qiu, and Huang 2017; Alice et al. 2019) adopting the insightful idea of domain adversarial training have been proposed to extract domain-invariant representations across multiple domains. 2) Reserving Differences. These approaches  (Ma et al. 2018; Tang et al. 2020) adopt multi-branch network architectures with domain-specific towers, gates, and even experts, implementing domain-specific parameters to avoid domain conflict caused by dataset bias across domains. As shown in Figure 2, these methods are becoming more and more complex with sophisticated network architectures or loss functions.
@@
Figure 1: (a)-(b): Review examples from a recommender system benchmark named Amazon with various styles and keywords (shown by different colors), as well as visual examples from a computer vision dataset named Office-Home with various appearances and backgrounds, reveal the main challenge of dataset bias. (c)-(d): The distribution of sample number across domains is naturally imbalanced or even long-tailed, indicating another major challenge of domain domination.
Further, different from multi-task learning which focuses on tackling different tasks within a single domain, multi-domain learning shares the same label space across multiple domains with different marginal distributions. We thus calculated the distribution of sample amount across domains on several benchmark datasets and the results are shown in Figure 1. In practice, the distributions of sample amount across domains are usually twisted, imbalanced, or even long-tailed, causing another main challenge of multi-domain learning: domain domination. In this case, some tailed domains may have much fewer examples and it would be difficult to train satisfied models on them. Without proper network designs or loss functions, the model will be easily dominated by the head domains with many more examples and shift away from the tailed ones with few examples.
Realizing the main challenges of dataset bias and domain domination in multi-domain learning, we aim to propose a general, effective, and efficient method to tackle these obstacles all at once. We first rethought the development of multi-domain learning and found that the approaches in this field are becoming more and more sophisticated, consisting of multifarious network architectures or complex loss functions with many trade-off hyperparameters. By assigning different parameters across domains, these designs may be beneficial in some cases but they will include more parameters and enlarge computation cost, as well as introduce much more hyperparameters. For example, the latest state-of-the-art method named PLE has to tune the numbers of domain-specific experts and domain-agnostic experts in each layer and design the network structures of each expert, each tower, and each gate network.
Motivated by the famous quote of Albert Einstein, “everything should be made as simple as possible, but no simpler”, we proposed a frustratingly easy multi-domain learning method named Decoupled Training (D-Train). D-Train is a training strategy based on the original frustratingly easy but most general shared-bottom architecture. D-Train is a tri-phase general-to-specific training strategy that first pre-trains on all domains to warm up a root model, then post-trains on each domain by splitting into multi-heads, and finally fine-tunes the heads by fixing the backbone, enabling decouple training to achieve domain independence. Despite its extraordinary simplicity and efficiency, D-Train performs remarkably well in extensive evaluations of various datasets from standard benchmarks to applications of satellite imagery and recommender systems. In summary, this paper has the following contributions:
@@
To tackle the dataset bias problem in multi-domain learning, various approaches (Liu, Qiu, and Huang 2017; Alice et al. 2019) have been proposed from the perspective of domain alignment, by adopting the insightful idea of domain adversarial training to extract domain-invariant representations across domains. There are mainly two categories of domain adaptation formulas: covariate shift (Quionero-Candela et al. 2009; Pan and Yang 2010; Long et al. 2015; Ganin and Lempitsky 2015) and label shift (Lipton, Wang, and Smola 2018; Azizzadenesheli et al. 2019; Alexandari, Kundaje, and Shrikumar 2020), while we focus on the former in this paper since it is more relevant with the topic of MDL. Recent deep domain adaptation methods tackle domain shifts from the perspectives of either moment matching or adversarial training, in which the former aligns feature distributions by minimizing the distribution discrepancy across domains (Long et al. 2015; Tzeng et al. 2014; Long et al. 2017). Further, domain adversarial neural network (DANN) (Ganin et al. 2016) becomes the mainstream method in domain adaptation. It introduces a domain discriminator to distinguish the source features from the target ones, while the feature extractor is designed to confuse the domain discriminator. In this way, the domain discriminator and feature extractor are competing in a two-player minimax game. Its natural extension to MDL is DANN-MDL. Later, CDAN (Long et al. 2018) further tailors the discriminative information conveyed in the classifier predictions into the input of the domain discriminator, whose natural extension to MDL is CDAN-MDL. Following the main idea of the minimax game, several variants of adversarial training methods (Pei et al. 2018; Tzeng et al. 2017; Saito et al. 2018; Zhang et al. 2019) were proposed. MulANN (Alice et al. 2019) tailors the insight of domain adversarial training into the MDL problem by introducing a domain discriminator into the shared model with a single head. In contrast, ASP-MTL (Liu, Qiu, and Huang 2017) includes a shared-private model and a domain discriminator.
@@
Another series of methods (Ma et al. 2018; Tang et al. 2020; Sheng et al. 2021) for multi-domain learning adopt multi-branch network architectures and develop domain-specific parameters (Dredze, Kulesza, and Crammer 2010) to avoid domain conflict caused by dataset bias across domains. Among them, Shared Bottom (SB) (Ruder 2017) is the frustratingly easy but effective one. Further, MoE (Jacobs et al. 1991) and its extension of MMoE (Ma et al. 2018) adopt the insightful idea of the mixture of experts to learn different mixture patterns of experts assembling, respectively. PLE (Tang et al. 2020) explicitly separates domain-shared and domain-specific experts to alleviate harmful parameter interference across domains. Note that, PLE further applies progressive separation routing with several deeper layers but we only adopt one layer for a fair comparison with other baselines. Other MDL methods focus on maintaining shared and domain-specific parameters by confidence-weighted combination (Dredze, Kulesza, and Crammer 2010), domain-guided dropout (Xiao et al. 2016), or prior knowledge about domain semantic relationships (Yang and Hospedales 2015). Meanwhile, various task-specific MDL approaches have been proposed for computer vision (Rebuffi, Bilen, and Vedaldi 2018; Mancini et al. 2020; Nam and Han 2016; Rebuffi, Bilen, and Vedaldi 2017; Li and Vasconcelos 2019; Fourure et al. 2017), natural language processing (Wu and Guo 2020; Williams 2013; Pham, Crego, and Yvon 2021) and recommender system (Hao et al. 2021; Chen et al. 2020; Du et al. 2019; Gu et al. 2021; Li et al. 2021b, a; Salah, Tran, and Lauw 2021). The comparison between these MDL methods is summarized in Figure 2.
@@
This paper aims to propose a simple and effective method for MDL. Given data points from multiple domains $\{\mathcal{D}_{1},\mathcal{D}_{2},...,\mathcal{D}_{T}\}$, $T$ is the domain number. Denote $\mathcal{D}_{t}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n_{t}}$ the data from domain $t$, where $\mathbf{x}_{i}$ is an example, $\mathbf{y}_{i}$ is the associated label and $n_{t}$ is the sample number of domain $t$. Denote the shared backbone $\psi$ and the domain-specific heads $\{{h}_{1},{h}_{2},...,{h}_{T}\}$ respectively. The goal of MDL is to improve the performance of each domain $\mathcal{D}_{t}$.
As mentioned above, the proposed Decoupled Training (D-Train) is a frustratingly easy multi-domain learning method based on the original shared-bottom architecture. D-Train takes a general-to-specific training strategy. It first pre-trains on all domains to warm up a root model. Then, it post-trains on each domain by splitting into multi-heads. Finally, it fine-tunes the heads by fixing the backbone, enabling decouple training to achieve domain independence.
@@
As mentioned before, jointly training a single model obscures domain distinctions, leading to domain conflict caused by the specificity of different domains. To reflect the domain specificity and achieve satisfactory performance for each domain, we adopt the shared-bottom architecture that has a shared feature extractor $\psi$ and various domain-specific heads $\{{h}_{1},{h}_{2},...,{h}_{T}\}$ to tackle the challenge of dataset bias across domains. With this design, the parameters of the feature extractor will be updated simultaneously by gradients of samples from all domains, but the parameters of domain-specific heads are trained on each domain. The optimization function of the post-training phase is formalized as
@@
Fine-tune: Decouple-Train for Independence
Regarding the benefits of domain-specific parameters across domains, a natural question arises: Can the shared-bottom architecture fully solve the problem of domain domination? To answer this question, we adopt the Euclidean norm to calculate the parameter update between the domain-specific heads $\{{h}_{1},{h}_{2},...,{h}_{T}\}$ and the domain-agnostic head ${h}_{0}$ at the pre-training phase on DomainNet. As revealed in Table 2, domain Real has more examples than other domains and is believed to dominate the training process. Our analysis in Figure 4(a) also verified that domain Real ($h_{3}$) is far away from the initial head. Since the parameter update at each training step $i$ is $h_{t}^{(i+1)}=h_{t}^{(i)}-\eta\frac{\partial L}{\partial h_{t}^{(i)}}$, we can easily attain the cumulative parameter update when the training converges: $h_{t}^{(T)}=h_{t}^{(0)}-\eta\sum_{i=0}^{T-1}\frac{\partial L}{\partial h_{t}^{(i)}}$, by repeating the calculation as $h_{t}^{(1)}=h_{t}^{(0)}-\eta\frac{\partial L}{\partial h_{t}^{(0)}},h_{t}^{(2)}=h_{t}^{(1)}-\eta\frac{\partial L}{\partial h_{t}^{(1)}},\cdots,h_{t}^{(T)}=h_{t}^{(T-1)}-\eta\frac{\partial L}{\partial h_{t}^{(T-1)}}$. Here, domain Real ($h_{3}$) has the largest value of the cumulative parameter update, we can say that it still dominates the training process. Hence, though the challenge of dataset bias across domains can be alleviated by introducing domain-specific heads in the shared-bottom architecture, the domain domination problem still exists after the post-train phase. This is reasonable since the domain-shared parameters of the backbone will be dominated by the head domains. To this end, we propose a decoupling training strategy by fully fixing the parameters of the feature extractor to achieve domain independence. Formally,
In this way, the parameters of the domain-specific heads will be learned by samples from each domain. With this kind of domain-independent training, the head domains will no longer dominate the training of the tailed domains at this phase. Further, the parameter update between phases becomes more balanced across domains as shown in Figure 4(b). Meanwhile, the training curves as shown in Figure 3(b) reveal that fine-tuning across domains further improves performance.
(a) After Post-train
(b) After Fine-tune
@@
Why Does D-Train Work?
As mentioned before, dataset bias and domain domination are the main challenges of multi-domain learning. All of the existing works, including the methods from the perspectives of both seeking commonalities or reserving differences, have to face the challenge of seesaw effect. With the shared parameters across domains, these methods will influence each other. Specifically, as shown in Figure 3(b),different domains achieve the optimal performance at different time stamps. When the problem of domain domination or domain conflict caused by dataset bias cannot be ignored, the MDL model will struggle to find an optimal solution for all domains. On the contrary, with the proposed decoupling training strategy, different domains in D-Train will train independently at the fine-tuning phase, enabling it the flexibility to achieve an optimal solution for each domain at different time stamps.
Further, we visualized the decision boundaries of different domains on Two-Moon with different scales and distributions in Figure 5. Caused by dataset bias and domain domination, the decision boundary of existing methods like Shared Bottom have some conflict areas, while D-Train can further adjust them by domain-independent fine-tuning.
(a) Shared Bottom
(b) D-Train
Figure 5: Decision boundaries on Two-Moon, where numbers in the legend indicate the accuracy of each domain.
@@
In this section, we compared the proposed D-Train method with baselines. Note that, domains are balanced by setting a fixed ‘batch_size’ for every domain, i.e., the ‘batch_size’ for each domain is $20$ if there are $80$ examples for $4$ domains in a mini-batch. We adopted this balanced version for all baselines and the proposed method because we found that the balanced implementation works much better than the original imbalanced one, e.g., Shared Bottom achieves an average accuracy of $75.9$ over $74.1$ in two versions.
@@
In this section, we adopt two standard benchmarks in the field of domain adaptation with various dataset scales, where Office-Home is in a low-data regime and DomainNet is a large-scale one.
@@
Office-Home is a standard multi-domain learning dataset  (Venkateswara et al. 2017) with $65$ classes and $15,500$ images from four significantly different domains: Art, Clipart, Product, and Real-World. As shown in Figure 1(b), there exist challenges of dataset bias and domain domination in this dataset. Following existing works on this dataset, we adopt ResNet-50 as the backbone and randomly initialize fully connected layers as heads. We set the learning rate as $0.0003$ and batch size as $24$ in each domain for D-Train and all baselines.
As shown in Table 1, Separately Training is a strong baseline and even outperforms Jointly Training, since the latter obscures domain distinctions and cannot tackle the dataset bias across domains. Multi-domain learning methods from the perspective of domain alignment work much better by introducing a domain discriminator and exploiting the domain information. However, applying domain alignment is not an optimal solution in MDL since the domain gap can only be reduced but not removed. Finally, the proposed D-Train consistently improves on all domains, even the tailed domain of “Art”. D-Train achieves a new state-of-the-art result with an average accuracy of $86.0$.
@@
DomainNet (Peng et al. 2019) is a large-scale multi-domain learning and domain adaptation dataset with $345$ categories. We utilize $4$ domains with different appearances including Clipart, Painting, Real, and Sketch where each domain has about $40,000$ to $200,000$ images. Following the code base in the Transfer Learning Library, we adopt mini-batch SGD with the momentum of $0.9$ as an optimizer, and the initial learning rate is set as $0.01$ with an annealing strategy. We adopt ResNet-101 as the backbone since DomainNet is much larger and more difficult than the previous Office-Home dataset. Meanwhile, the batch size is set as $20$ in each domain here for D-Train and all baselines.
As shown in Table 2, the proposed D-Train outperforms all baselines, no matter measured by average accuracy or worst accuracy in all domains. Since the domain of “Real” has much more samples than other domains, it achieves competitive performance while separately training. However, other domains benefit from training via D-Train.
@@
In this section, we adopt a popular satellite dataset named Functional Map of the World (FMoW)  (Christie et al. 2018), which aims to predict the functional purpose of buildings and land use on this planet, contains large-scale satellite images with different appearances and styles from various regions: Africa, Americas, Asia, Europe, and Oceania. FMoW is a natural dataset for multi-domain learning. D-Train and all baselines in this section are implemented in a popular open-sourced library named WILDS since it enables easy manipulation of this dataset. Each input $\mathbf{x}$ in FMoW is an RGB satellite image that is resized to $224\times 224$ pixels and the label $y$ is one of $62$ building or land use categories. For all experiments, we follow (Christie et al. 2018) and use a DenseNet-121 model (Huang et al. 2017) pre-trained on ImageNet. We set the batch size to be $64$ on all domains. Following WILDS, we report the average accuracy and worst-region accuracy in all multi-domain learning methods.
@@
As shown in Table 3, it’s not wise to train a separate model for each domain on FMoW, since the data on some domains is extremely scarce. FMoW is a difficult and scarce dataset, e.g., Oceania has only 15k images, causing some categories to involve only a dozen examples. Joint Train is also not optimal, because conflicts widely exist in some domains. Note that D-Train also outperforms all baselines in this difficult dataset.
@@
Table 4: AUC (%) on Amazon Product Review for multi-domain learning.